40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3362     Entry time: Wed Aug 4 20:15:07 2010
Author: Jenne 
Type: Update 
Category: PEM 
Subject: Yay! Guralps work again! 

After much hassle, the Guralp cable from the ADC Out of the breakout box to the ADC is fixed, and everything is plugged in and working again.  The seismometers are back in their regular positions at the ends of the MC, ready for some excellent seis/MC combo data. 

I solidified the change of putting the Gur2Z channel into a different BNC input on the ADC.  The C1ADCU_PEM.ini file has been changed so that what used to be the Ranger's channel is now recognized as Gur2Z. 

Also, I changed the same .ini file to reflect Koji's move of ACC_MC1_Z to the old AUDIO_MIC2 channel, so now all 6 Accelerometer channels have the same calibrations again.

Another big change is the change from old-left-handed convention to new-right-handed convention.  The seismometers are aligned the same way they always have been (with the North-South markers aligned with the MC), but now the North-South output is plugged into the BNC on the ADC that is associated with Gur*_X, and the East-West output is plugged into the ADC channel associated with Gur*_Y.  This is true for both Guralp Seismometers. 

So, now we have:

Gur1_X = Gur1_NS = ADC#10

Gur1_Y = Gur1_EW = ADC#11

Gur1_Z = Gur1_Vert = ADC#12

Gur2_X = Gur2_NS = ADC#2

Gur2_Y = Gur2_EW = ADC#3

Gur2_Z = Gur2_Vert = ADC#24

SEIS_Ranger_Y = no longer in the .ini file

ELOG V3.1.3-